# GiBy: A Giant-Step Baby-Step Classifier For Anomaly Detection In Industrial Control Systems

### v1

@@ VIII Conclusions and Future Work
We presented a simple yet powerful anomaly detector that trained four set of bounds — the min-max bounds for the giant-step, the rate of change bounds for baby-step, and the extended detection algorithm bounds, for both giant and baby-steps. The detector acted as a one-class classifier. Any tested value outside trained bounds were flagged as an anomaly. The transparent design of GiBy anomaly detector made it straightforward to implement and easy to interpret. The explainability provided pinpoints the sensor and actuation state for which anomaly was detected, and what bounds it breached. The experiments showed that testing for an anomaly and providing an explanation took around of a second per sensor. This made it useful for implementation in systems where near real-time decisions are made, or on devices that are resource constrained.
In our experiments, we observed in TABLE VIII that some attacks were well within the range of normal behavior and could not be detected. We emphasize that detecting all attacks in an attack table (TABLE VIII) is not the same as detecting all attacks on a system. Attack tables, such as those provided by SWaT are useful to check whether the detector detects different types of attacks, based on attacker capabilities. However, detecting some attacks are expected to remain difficult for all anomaly detectors, due to entropy loss in training and indistinguishability from normal operational data. Due to the undetectable attacks discussed in Section V-1, in future work we will explore other types of detection using digital signature verification and encryption to hide the inferences on sensor readings across OT, segregating and air-gapping high risk sensors and PLCs, paying closer attention to attack prevention and fast recovery strategies that do not compromise plant and personnel safety in the meantime.

### v3

@@ 8 Conclusions and Future Work
We presented a simple yet powerful anomaly detector that trained four set of bounds — the min-max bounds for the giant-step, the rate of change bounds for baby-step, and the extended detection algorithm bounds, for both giant and baby-steps. The detector acted as a one-class classifier. Any tested value outside trained bounds were flagged as an anomaly. The transparent design of GiBy anomaly detector made it straightforward to implement and easy to interpret. The explainability provided pinpoints the sensor and actuation state for which anomaly was detected, and what bounds it breached. The experiments showed that anomaly detection and explanation took around of a second per sensor (on average). This makes it useful for implementation in systems where near real-time decisions are made, or on devices that are resource constrained. Due to the undetectable attacks example discussed in Section 5.0.2, in future work we will explore other types of detection using digital signature verification and encryption to hide the inferences on sensor readings across OT, segregating and air-gapping high risk sensors and PLCs, paying closer attention to attack prevention and fast recovery strategies that do not compromise plant and personnel safety in the meantime.